OpenStudy (anonymous):

Solve the equation on the interval 0 ≤ θ < 2π. please help 3 cot^2 θ - 4 csc θ = 1 a)pi/6 , 5pi/6 b)pi/6 c)7pi/6 d)7pi/6, 11pi/6

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles \[3\cot^2\theta-4\csc\theta=1\\ \frac{3\cos^2\theta}{\sin^2\theta}-\frac{4}{\sin\theta}=1\\ 3\cos^2\theta-4\sin\theta=\sin^2\theta\\ 3\left(1-\sin^2\theta\right)-4\sin\theta=\sin^2\theta\\ 3-3\sin^2\theta-4\sin\theta-\sin^2\theta=0\\ 3-4\sin\theta-4\sin^2\theta=0\\ (3+2\sin\theta)(1-2\sin\theta)=0\] Solve the following: \[\begin{cases}3+2\sin\theta=0\\ 1-2\sin\theta=0\end{cases}\] in the given interval. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

okay thanks you

OpenStudy (anonymous):

yw